# Document-Level In-Context Few-Shot Relation Extraction via Pre-Trained Language Models

### v2

@@ 4.3 Computing Knowledge Triplet Probabilities
$p(s\,|\,C_{l},d_{i},r)$ =$\mathrm{len}(s)]{\prod_{k=1}^{\mathrm{len}(s)}p(s_{k}\,|\,s_{<k},C_{l},d_{i},r)},$ (4)
$p(o\,|\,s,C_{l},d_{i},r)$ =$\mathrm{len}(o)]{\prod_{k=1}^{\mathrm{len}(o)}p(o_{k}\,|\,o_{<k},s,C_{l},d_{i},r)},$ (5)
@@ 5 Experimental Setup
We evaluate our framework using the DocRED (Yao et al., 2019), the largest document-level relation extraction dataset publicly available. DocRED includes 96 relation types and comprises three sets: (1) a distantly-supervised set with 101,873 documents, (2) a human-annotated train set with 3053 documents, and (3) a human-annotated development set with 998 documents. We provide further details about the DocRED dataset in the Appendix B. Importantly, in our experimental setup, we use distantly-supervised set for in-context few-shot learning ($\mathcal{D}^{\text{dist}}$) and evaluate the performance on the development set. Thereby, we ensure that our framework is solely trained without human annotation. To better understand the performance of our REPLM, we thus later also perform additional experiments (see Sec. 7) using sentence-level relation extraction datasets.
@@
Baselines. We evaluate our framework against state-of-the-art methods for relation extraction that scale to document-level and, for comparability, do not require named entity recognition pipelines (see Table 1). These are: (1) REBEL (Cabot and Navigli, 2021) applies triplet linearization to extract the relations from the given document. (2) We also compare against REBEL-sent, which extracts the relations from a document in sentence-by-sentence manner. We include this variant because REBEL is originally trained at sentence level and, as we show later, it is the best-performing baseline method for the sentence-level relation extraction. Note that REBEL is the only baseline from the literature that does not require a named entity recognition pipeline for document-level relation extraction.
@@
In our dataset, we use REBEL-large from Hugging Face [^3] https://huggingface.co/Babelscape/rebel-large, which is pre-trained by a tailored REBEL dataset [^4] https://huggingface.co/datasets/Babelscape/rebel-dataset. We note that REBEL-large is further fine-tuned on the human-annotated training set of DocRED, which may give it an (unfair) advantage. Hyperparameter selection and early-stopping are done based on some portion of the development set. This is again an (unfair) advantage for REBEL, whereas our REPLM does not need such dev set.
@@
REPLM variants: We compare two variants of our proposed framework: (1) REPLM is the original variant as described above. Therein, we use fixed parameters. We also report a sensitivity analysis in Appendix J, where we vary the parameters to see that our performance is robust to different choices. (2) REPLM (params adj) is a variant for which the hyperparameters (e. g., temperature, threshold) are selected based on the training set.
@@
Ablations: We perform ablation studies to assess the contribution of different components in our framework. Our ablations are: (1) REPLM (random fixed) randomly selects a single set of $K$ documents for each relation [^5] We also explored finding the “best” documents (Appendix K) of a relation. It requires evaluation against human-annotation and still performs worse than REPLM. Hence, we exclude its results.. However, the set is fixed across all evaluations. (2) REPLM (random all) randomly selects a set of $K$ documents for each relation and for each evaluation. (3) REPLM (best $\ominus$) selects the top-$K$ documents for each relation and each development document according to the cosine similarity. In-context examples are ordered from most similar to the least similar. (4) REPLM (best $\oplus$) similarly selects top $K$ documents for each relation and each development document. This time, in-context examples are ordered in reverse order, from least similar to most similar. We include these two alternatives to evaluate the effect (if any) of recency bias (Lu et al., 2022a; Hongjin et al., 2023).
@@
Implementation. We use GPT-JT [^6] https://huggingface.co/togethercomputer/GPT-JT-6B-v1 ($\sim$6B parameters) as our pre-trained LM for in-context few-shot learning. Appendix E provides all details of our REPLM framework.
@@ 6.1 Overall Performance
Table 2: Document-level relation extraction results. Shown: Micro F1 scores.
@@
In our ablation study, we compare different variants of our REPLM to understand the source of performance gains (see Fig. 2 and Table 2). We make the following observations: (1) Retrieving the best in-context examples improves the performance compared to random examples by more than 48 % (REPLM (best $\ominus$) and REPLM (best $\oplus$) vs. REPLM (random fixed) and REPLM (random all)). (2) We do not observe that a recency bias plays a decisive role for our results, as both REPLM (best $\ominus$) and REPLM (best $\oplus$) reach a similar performance. (3) Our complete framework brings a significant improvement over REPLM (best $\ominus$) and REPLM (best $\oplus$) (+18 %) by aggregating multiple sets of most relevant in-context examples, thus establishing the importance of using multiple sets.
Insights. We conjecture that our REPLM extracts more relations than REBEL, as it further identifies the missing annotations in DocRED. In Appendix F, we empirically validate that, for each relation type, there are dev documents with no annotation, which are semantically similar to the documents containing at least one knowledge triplet of a relation. It suggests that these dev documents potentially include the relation but lack the annotation. To confirm this conjecture, we perform a manual validation that we examine cases where our REPLM fails. We find that many relations extracted by our method are actually correct but considered as false positives due to missing human annotations. For example, our REPLM generates the relation (author, Chaosmosis, Félix Guattari) but it is not annotated and therefore marked as incorrect (see right part of Fig. 1). We provide additional examples in the Appendix G.
@@ 6.2 Comparison Against External Knowledge
Figure 3: F1 scores per relation type (darker = better). Here, the comparison is made against annotations that additionally make use of external knowledge and should thus more closely reflect the ground-truth. Relation types are arranged in decreasing order of their number of knowledge triplets. For visibility, F1 scores are normalized by the maximum value for each relation.
@@
Ground-truth via external knowledge: To locate the missing annotations in DocRED, we aggregate all relations extracted from all methods on all documents. We then check the correctness of the extracted relations via an external KB. Specifically, we leverage the pipeline from HELM (Liang et al., 2022) and check if the generated knowledge triplet exists in Wikidata (Vrandečić and Krötzsch, 2014). We add all the matched triplets to the existing list of ground-truth triplets from DocRED and repeat the evaluation. As a result, total number of relations in the development set increased from 12,212 to 18,592.888We note that the increase in the number of relations does not necessarily imply an improvement in the F1 score for our REPLM. The extracted relations are still filtered by the probability threshold $\theta$, which, in turn, reduces the recall (and possibly the F1 score).
@@
Results: For our new dataset with external ground-truth, our framework outperforms REBEL by a considerable margin across most relation types (Fig. 3) [^9] The details of evaluation via external KB are in Appendix I.. This is also seen in the overall performance (Table 3). For example, our REPLM improves the F1 score over REBEL by more than 59 % (32.33 vs. 20.30). The improvement for REPLM (params adj) is even larger and amounts to 80 % (36.51 vs 20.30).
7 Ablation Study: Performance for Sentence-Level Relation Extraction
After showing the effectiveness of our REPLM in document-level relation extraction, we additionally perform sentence-level relation extraction as an ablation study, albeit not our primary benchmarking focus. Despite this, our REPLM still matches or surpasses the performance of fine-tuned LMs.
Datasets. We conduct extensive experiments on three sentence-level relation extraction benchmark datasets, namely CONLL04 (Roth and Yih, 2004), NYT (Riedel et al., 2010), and ADE (Gurulingappa et al., 2012). We provide the details of each dataset in Appendix C.
Baselines. We compare REPLM against fine-tuned LMs that achieve state-of-the-art results: SpERT (Eberts and Ulges, 2020), Table-sequence (Wang and Lu, 2020), TANL and TANL (multi-dataset) [^10] This version is jointly trained on all 3 datasets. (Paolini et al., 2021), REBEL (w/o pre-training) and REBEL (Cabot and Navigli, 2021) [^11] The same REBEL from Sec. 6 that is pre-trained by an external dataset before fine-tuning.
Performance. Table 4 shows the micro F1 scores of each method on CONLL04, NYT, and ADE datasets. Our REPLM achieves the best performance on ADE against all the supervised methods with an F1 score of 82.5. We observe that our REPLM performs on par with the state-of-the-art methods on CONLL04 dataset with an F1 score of 72.9, while REBEL attains the best score via an external dataset. For NYT, our REPLM achieves a lower score than state-of-the-art methods. The reason is, the entire NYT is curated via distant supervision, which results in noisy relations. As the baseline methods are fine-tuned with the same dataset, they can memorize the same knowledge triplets from the training set, and therefore, achieve spuriously high performance. The detailed explanation is in Appendix C.
The above analysis also justifies our choice of using REBEL, which is the best of the baselines. It is further the only model that can extract relations from the documents without named entity annotation. Hence, the above ablation also explains why we choose REBEL as the main baseline in Sec. 6.
Table 4: Sentence-level relation extraction. Shown: Micro F1.
Method CONLL04 NYT ADE
SpERT (Eberts and Ulges, 2020) 71.5 _ 79.2
Table-sequence (Wang and Lu, 2020) 73.6 _ 80.1
TANL (Paolini et al., 2021) 71.4 90.8 80.6
TANL (multi-dataset) (Paolini et al., 2021) 72.6 90.5 80.0
REBEL (w/o pre-training) (Cabot et al., 2021) 71.2 91.8 81.7
REBEL (Cabot et al., 2021) 75.4 92.0 82.2
REPLM (ours) 72.9 81.0 82.5
Higher is better. Best value in bold.
@@ 7 Ablation Study: Performance for Sentence-Level Relation Extraction
Further insights. What is the effect of the number of in-context examples? We repeat the same experiment on CONLL04 with a varying number of in-context examples. Fig. 3(a) shows the F1 score on each relation and the overall score by varying $K$ from 3 to 11. We observe that, in general, more in-context examples yield better F1 scores. Informed by this observation, we used the highest number of in-context examples that fit into the context window for our experiments, which is $K=5$ for document-level relation extraction. Detailed results are given in Appendix L.
@@
Is REPLM actually learning to extract relations? Or does it only retrieve facts from memory? We design a novel experiment to identify whether our REPLM is learning to extract relations from the input text or it is simply retrieving the facts from its memory. To the best of our knowledge, we are the first to shed more light on the models’ learning ability for the relation extraction task. For this experiment, we replaced all the entities with random names in CONLL04 dataset (for both training and test set) that are not mentioned anywhere in the web. Fig. 3(b) compares the performance against the original dataset. The overall performance decreases only slightly when using the random entities (F1 score of 70.47 vs. 72.9), which is still on par with the state-of-the-art. Therefore, it confirms that our REPLM is an effective method in learning to extract the relations from the context. We provide the experiment details and elaborate on the reasons of slight performance decrease in Appendix M.

### v4

@@ 4.3 Computing Knowledge Triplet Probabilities
[…]
@@ 5 Experimental Setup
We perform an extensive evaluation of our framework using the DocRED (Yao et al., 2019), the largest document-level relation extraction dataset publicly available. DocRED includes 96 relation types and comprises three sets: (1) a distantly-supervised set with 101,873 documents, (2) a human-annotated train set with 3053 documents, and (3) a human-annotated dev set with 998 documents. We provide further details about the DocRED dataset in Appendix B.1. Importantly, in our experimental setup, we use a distantly-supervised set for in-context few-shot learning ($\mathcal{D}^{\text{dist}}$) and evaluate the performance on the development set. Thereby, we ensure that our framework is solely trained without human annotation. To better understand the performance of our REPLM, we thus later also perform additional experiments (see Sec. 7) using sentence-level relation extraction datasets.
@@
Baselines. We evaluate our framework against state-of-the-art methods for relation extraction that scale to document-level and, for comparability, do not require named entity recognition pipelines (see Table 1). These are: (1) REBEL (Cabot & Navigli, 2021), applying triplet linearization to extract relations from the document. (2) REBEL-sent, extracting relations in a sentence-by-sentence manner. We include this variant because REBEL is originally trained at sentence level and, as shown later, is the best-performing baseline for sentence-level relation extraction. Note that REBEL is the only baseline from the literature not requiring a named entity recognition pipeline for document-level relation extraction.
@@
In our experiments, we use REBEL-large from Hugging Face [^3] https://huggingface.co/Babelscape/rebel-large, which is pre-trained by a tailored REBEL dataset [^4] https://huggingface.co/datasets/Babelscape/rebel-dataset. We note that REBEL-large is further fine-tuned on the human-annotated training set of DocRED, which may give it an (unfair) advantage. Hyperparameter selection and early stopping are based on the development set, which is again, an advantage not needed by our framework.
@@
REPLM variants: We compare two variants of our framework: (1) REPLM is the original variant as described above. Therein, we use fixed parameters. A sensitivity analysis in Appendix J shows that the performance robustness to different parameter choices. (2) REPLM (params adj) is a variant for which the hyperparameters (e. g., temperature, threshold) are selected based on the training set.
@@
We assess the contribution of different components in our framework. For this, we run an extensive series of experiments using the following variants: (1) REPLM (random fixed) randomly selects a single set of $K$ documents for each relation [^5] We also explored finding the “best” documents (Appendix K) of a relation. It requires evaluation against human-annotation and still performs worse than REPLM. Hence, we exclude its results.. However, the set is fixed across all evaluations. (2) REPLM (random all) randomly selects a set of $K$ documents for each relation and for each evaluation. (3) REPLM (best $\ominus$) selects the top-$K$ documents for each relation and each document according to the cosine similarity. In-context examples are ordered from most similar to the least similar. (4) REPLM (best $\oplus$) similarly selects top $K$ documents for each relation and each document. This time, in-context examples are ordered in reverse order, from least similar to most similar. We include these two alternatives to evaluate the effect (if any) of recency bias (Hongjin et al., 2023, Lu et al., 2022a).
@@
Implementation. We mainly use GPT-JT [^6] https://huggingface.co/togethercomputer/GPT-JT-6B-v1 ($\sim$6B parameters) as our pre-trained LM for in-context few-shot learning. Our additional experiments (Sec. 7 and Sec. 8) show that other LMs can be seamlessly incorporated into our REPLM, such as OpenAI’s GPT models or Meta’s Llama models. Appendix E provides all details of our framework.
@@ 6.1 Overall Performance
Table 2: Document-level relation extraction results. Shown: Micro F1.
@@
We further compare different variants of our REPLM to understand the source of performance gains (see Fig. 2 and Table 2). (1) Retrieving the best in-context examples improves the performance compared to random examples by more than 48 % (REPLM (best $\ominus$) and REPLM (best $\oplus$) vs. REPLM (random fixed) and REPLM (random all)). (2) We do not observe that a recency bias plays a decisive role in our results, as both REPLM (best $\ominus$) and REPLM (best $\oplus$) reach a similar performance. (3) Our complete framework brings a significant improvement over REPLM (best $\ominus$) and REPLM (best $\oplus$) (+18 %) by aggregating multiple sets of most relevant in-context examples, thus establishing the importance of using multiple sets.
Insights. We conjecture that our REPLM extracts more relations than REBEL, as it further identifies missing annotations in DocRED. In Appendix F, we empirically validate that, for each relation type, some dev documents have no annotation but are semantically similar to those containing at least one knowledge triplet. This suggests these dev documents include the relation but lack the annotation. To confirm, we manually validate cases where our REPLM fails. We find many relations extracted by our method are correct but considered false positives due to missing annotations. For example, REPLM generates the relation (author, Chaosmosis, Félix Guattari) but it is not annotated and thus marked as incorrect (see right part of Fig. 1). Additional examples are in Appendix G.
@@ 6.2 Comparison Against External Knowledge
Ground-truth via external knowledge: To locate missing annotations in DocRED, we aggregate all relations extracted from all methods on all documents. We then check the correctness of the extracted relations via an external KB. Specifically, we leverage the pipeline from HELM (Liang et al., 2022) and check if generated knowledge triplets exist in Wikidata (Vrandečić & Krötzsch, 2014). We add all matched triplets to the existing list of ground-truth triplets from DocRED and repeat the evaluation. As a result, total number of relations in development set increased from 12,212 to 18,592.888We note that the increase in the number of relations does not necessarily imply an improvement in the F1 score for our REPLM. The extracted relations are still filtered by the probability threshold $\theta$, which, in turn, reduces the recall (and possibly the F1 score).
@@
Results: For DocRED with external ground-truth, our framework outperforms REBEL by a considerable margin across most relation types (Fig. LABEL:fig:res_3) [^9] The details of evaluation via external KB are in Appendix I. and in the overall performance (Table 3). For example, our REPLM improves F1 score over REBEL by more than 59 % (32.33 vs. 20.30). The improvement for REPLM (params adj) is even larger and amounts to 80 % (36.51 vs 20.30).
7 Extensive Benchmarking across Additional Datasets
After showing the effectiveness of our complete REPLM in the largest available document-level relation extraction dataset, we now turn to both smaller document-level datasets and sentence-level datasets, and then now conduct one of the most extensive benchmarking studies in relation extraction. Specifically, we implement our framework with 5 different LLM backbones, and compare them across 6 relation extraction datasets against more than 30 baseline methods. Yet, unlike our framework, the baselines have large computational overhead (e.g., additionally requiring both named-entity recognition and fine-tuning).
REPLM variants: The five different LLM backbones of our framework are: GPT-JT, Llama-3.1-8B, Llama-3.1-70B, GPT-3.5-Turbo, and GPT-4o. [^10] For Llama, see Dubey et al. (2024). For GPT, see https://platform.openai.com/docs/models. Datasets: On top of DocRED, we consider two additional document-level relation extraction datasets: CDR (Li et al., 2016b) and GDA (Wu et al., 2019). [^11] Details about CDR and GDA are given in Appendix B.2. We consider three sentence-level datasets: CONLL04 (Roth & Yih, 2004), NYT (Riedel et al., 2010), and ADE (Gurulingappa et al., 2012). Baseline methods: All +30 baseline methods are only listed in Table 4 due to space.
Performance: Table 4 presents the micro-F1 scores for all methods across datasets. (1) Baseline methods show a clear divide: some can only be used for documents, while others only be used for sentences. This is due to that (a) sentence-level methods classify all named-entity pairs, which does note scale to documents, and that (b) document-level methods model inter- and intra-sentence relations explicitly, which does not apply to sentence-level datasets. In contrast, our framework handles all datasets and scales easily to larger models. (2) Adopting newer, stronger language models significantly boosts performance. For instance, on DocRED, the F1 score increased by $+$24.57 from GPT-JT to GPT-3.5-Turbo, and by $+$8.69 from GPT-3.5-Turbo to GPT-4o. This trend is consistent across datasets and between Llama-3.1-8B and Llama-3.1-70B. (3) Our REPLM with GPT-4o achieves the best performance on DocRED, CoNLL04, and ADE, and near-best results on CDR and NYT.
We further investigated cases where the best variant of our framework, i.e., REPLM (GPT-4o), did not achieve top performance. The issues stem from missing or inconsistent entity annotations in the biomedical datasets CDR and GDA. Baseline methods, trained with these annotations, implicitly overfit to them and avoid the issue. For example, our framework correctly identifies the triplet “Gene-Disease Association, complement receptor 1, insulin-dependent diabetes mellitus” in GDA, but it is marked as a false positive since “complement receptor 1” is only annotated as “CR1”, “C3bR”, and “CD35”. In NYT, noisy relations from the distantly-supervised dataset curation lead baseline methods to memorize triplets from training, thus inflating their performance incorrectly. Detailed analyses are provided in Appendix C.
Table 4: Evaluation of REPLM variants across datasets. Shown: Micro F1.
Document-level Sentence-level
Method DocRED CDR GDA CONLL04 NYT ADE
Document-level SciBERT (Beltagy et al., 2019) 65.1 82.5
methods LSR (Nan et al., 2020) 64.8 82.2
DHG (Zhang et al., 2020) 65.9 83.1
GAIN (Zeng et al., 2020) 61.22
JEREX (Eberts & Ulges, 2021) 40.41
HeterGSAN (Xu et al., 2021c) 60.18
DRN (Xu et al., 2021b) 61.39
SIRE Zeng et al. (2021) 61.60
SSAN (Xu et al., 2021a) 65.69 68.7 83.7
ATLOP (Zhou et al., 2021) 63.40 69.4 83.9
E2GRE (Huang et al., 2021) 58.72
DocuNet (Zhang et al., 2021) 64.55 76.3 85.3
EIDER (Xie et al., 2022) 64.79
SAIS (Xiao et al., 2022) 65.17 79.0 87.1
DREEAM (Ma et al., 2023) 67.41
DocRE-CLiP (Jain et al., 2024) 68.13
Sentence-level Neural Joint (Li et al., 2016a) 63.40
methods SpERT Eberts & Ulges (2020) 71.54 79.22
Table-sequence Wang & Lu (2020) 73.58 80.07
BILSTM + Att (Geng et al., 2020) 71.39
TANL Paolini et al. (2021) 71.48 90.83 80.61
TriMF (Shen et al., 2021) 72.35
CMAN (Zhao et al., 2021a) 72.97 81.14
CL (Theodoropoulos et al., 2021) 79.97
PFN (Yan et al., 2021) 83.20
REBEL Cabot & Navigli (2021) 27.52 75.41 92.02 82.23
TabERT (Ma et al., 2022) 72.60
BL (Ji et al., 2022) 72.62 81.33
STER (Zhao et al., 2022) 74.44 81.24
FedJ (Wang et al., 2023a) 72.35 82.37
PREFER (Liu et al., 2023) 75.66 84.98
GPT-RE Wan et al. (2023) 45.84
CodeIE Li et al. (2023) 53.12 32.22
Our framework REPLM (GPT-JT) 35.09 55.98 66.92 72.94 81.03 82.54
REPLM (Llama-3.1-8B) 55.50 67.36 71.07 69.43 84.55 87.11
REPLM (Llama-3.1-70B) 62.31 73.29 74.10 72.00 89.33 91.45
REPLM (GPT-3.5) 59.66 65.19 72.54 80.19 86.16 84.29
REPLM (GPT-4o) 68.35 73.62 74.11 85.22 90.12 92.17
Best values are in bold. CodeIE is the only baseline that does not require any model training.
8 Ablation Study
Table 5: Ablation of REPLM variants. Shown: micro F1.
Document-level Sentence-level
Backbone Variant DocRED CDR GDA CONLL04 NYT ADE
GPT-JT random context 21.14 39.21 57.48 62.48 58.92 76.37
best context 31.31 47.37 63.02 68.16 77.73 79.82
complete framework 35.09 55.98 66.92 72.94 81.03 82.54
GPT-3.5 random context 46.83 58.57 66.41 72.12 68.09 78.36
best context 54.31 62.16 68.40 74.09 83.81 81.63
complete framework 59.66 65.19 72.54 80.19 86.16 84.29
GPT-4o random context 52.29 67.41 69.30 77.19 73.09 85.55
best context 61.78 71.66 71.59 79.90 86.81 90.18
complete framework 67.47 73.62 74.11 85.22 90.12 92.17
Llama-3.1-8B random context 30.01 47.87 66.25 40.15 47.67 72.34
best context 40.85 50.38 68.46 53.03 62.51 81.18
complete framework 55.50 67.36 71.07 69.43 84.55 87.11
Llama-3.1-70B random context 52.02 63.63 68.99 34.54 60.91 82.37
best context 57.28 67.44 70.64 53.63 83.72 82.58
complete framework 62.31 73.29 74.10 72.00 89.33 91.45
Best values in bold. The std. dev. of random contexts are all below than 0.1 and thus
omitted for brevity.
Are the performance gains of REPLM robust across different datasets and different LM backbones? We present an ablation study demonstrating the effectiveness of our complete framework on six relation extraction datasets, evaluated across different LM backbones and variants of our own framework (i.e., random context vs. best context vs. complete framework). Our complete framework consistently outperforms retrieving only the best context, which, in turn, performs better than a random context. This pattern holds across both document-level and sentence-level datasets and all five backbone models. These findings demonstrate an important implication: whenever more powerful LMs become available, one can integrate them into our REPLM in a seamless manner and thereby achieve important performance gains for relation extraction tasks.
@@ 8 Ablation Study
What is the effect of the number of in-context examples? We repeat the same experiment on CONLL04 when varying the number of in-context examples ($K$). Fig. 3(a) shows (i) the F1 score for each relation and (ii) the overall score when varying $K$ from 3 to 11. We observe that, in general, more in-context examples yield better F1 scores. Informed by this observation, we used the highest number of in-context examples that fit into the context window for our main experiments, which is $K=5$ for document-level relation extraction. Detailed results are given in Appendix L.
@@
Is REPLM actually learning to extract relations? Or does it only retrieve facts from memory? We design a novel experiment to identify whether our REPLM is learning to extract relations from the input text or it is simply retrieving the facts from its memory. To the best of our knowledge, we are the first to shed more light on the models’ learning ability for the relation extraction task. For this experiment, we replaced all the entities with random names in CONLL04 dataset (for both training and test set) that are not mentioned anywhere on the web. Fig. 3(b) compares the performance against the original dataset. The overall performance decreases only slightly when using the random entities (F1 score of 70.47 vs. 72.9), which is still on par with the state-of-the-art. Therefore, it confirms that our REPLM is an effective method for learning to extract the relations from the context. We provide the experiment details and elaborate on the reasons of the slight performance decrease in Appendix M.